### v1

# 2]Dept. of Physics and Astronomy and LENS, Florence Univ., via Sansone 1, I-50019 Sesto Fiorentino, Italy

Abstract
Quantum machine learning is a new research field combining quantum information science and machine learning. Quantum computing technologies seem to be particularly well suited to solving problems in the health sector in an efficient way, because they may deal with large datasets more efficiently than classical AI.
Alzheimer’s disease is a neurodegenerative brain disorder that mostly affects elderly people, causing important cognitive impairments. It is the most common cause of dementia and it has an effect on memory, thought, learning abilities and movement control. This type of disease has no cure, consequently an early diagnosis is fundamental for reducing its impact. The analysis of handwriting can be effective for diagnosing, as many researches have conjectured. The DARWIN (Diagnosis AlzheimeR WIth haNdwriting) dataset contains handwriting samples from people affected by Alzheimer’s disease and a group of healthy people. Here we apply quantum AI to this use-case. In particular, we use this dataset to test kernel methods for classification task and compare their performances with the ones obtained via quantum machine learning methods. We find that quantum and classical algorithms achieve similar performances and in some cases quantum methods perform even better.
Our results pave the way for future new quantum machine learning applications in early-screening diagnostics in the healthcare domain.
keywords:
Quantum machine learning, Alzheimer’s disease, supervised learning, data classification, parametrized quantum circuit
Figure 1: The data obtained from the handwriting tasks are processed and mapped into quantum states through a parametrized quantum circuit (PQC). Then a classification algorithm is applied to distinguish the samples between ‘patient’ (P) and ‘healthy’ (H).
1 Introduction
Neurodegenerative diseases are incurable conditions caused by the progressive degeneration of nerve cells [1]. Alzheimer’s disease (AD) is one of the most common amongst them [2]. The predominant symptom in the early phase of AD is episodic memory impairment followed by progressive amnesia, a result of widespread brain damage. Since there is no cure for AD, it is critical to improve the approaches now used for diagnosis [3]. Handwriting is a task that involves both cognitive and motor functions to plan and properly execute the movements that are required with high coordination [4]. The analysis of handwriting and drawing dynamics is a valid non-invasive way for evaluating AD progression. For this purpose a protocol of 25 tasks was defined in Ref. [5]. Each handwriting/drawing task is described by 18 features and the data collected amongst a group of healthy people and patients made up the DARWIN (Diagnosis AlzheimeR WIth haNdwriting) dataset. In Ref. [6] it is described how this dataset can be exploited by machine learning models to discriminate between AD patients and healthy people.
Machine learning (ML) is the field of artificial intelligence (AI) that studies the development of learning models to recognize, for example, patterns in large sets of data [7]. These models can learn how to attribute outputs to unseen inputs for classification or regression tasks (supervised learning), or find a structure in the inputs (unsupervised learning) for clustering or dimensionality reduction [8]. A lot of work has been done in the past years to improve the performance of ML algorithms and make them capable of handling efficiently big datasets. One more recent and promising solution could be to use quantum machine learning (QML) techniques. In QML inputs are encoded in quantum states and computation is done on a quantum computer [9, 10]. Although quantum processing units (QPUs) are still subjected to relatively large noise rates, it seems that applications in ML can lead to remarkable results. For example, in Ref. [11] it is shown how variational quantum models, namely models based on updating parameters in a quantum circuit through classical optimization, can exhibit a significant learning advantage in solving a regression task with input data from the fashion-MNIST dataset. A review of the contributions of QML in medical image analysis is given in Ref. [12]. A previous study [13] approaches via a so-called variational quantum classifier (VQC) the problem of classification defined by the same dataset we consider in our work.
Here we focus in particular on quantum kernel methods, i.e. kernel methods in which the feature space is a space of quantum states. We use these algorithms to classify the samples and compare their performances with the ones achieved by classical algorithms. The paper is organized as follows. In Section 2 we describe how the DARWIN dataset was created and we define the classical and quantum kernel methods for classification. Then in Section 3 and 4 we report their performances, we discuss the results and draw our conclusions.
2 Methods
2.1 Dataset
The DARWIN dataset is composed of handwriting data collected according to the protocol defined in Ref. [5]. It includes 25 tasks that can be divided in the following categories: graphic tasks, like joining points and drawing geometrical figures; copy tasks, testing the ability of repeating more complex gestures to write letters, numbers and words; memory tasks, involving writing previously memorized words; and dictation tasks. The protocol was submitted to 174 participants: 89 AD patients and 85 healthy people, recruited so that the two groups would match in terms of age, education and gender. None of the participants was taking medications that influenced their cognitive abilities. The tasks were performed on paper sheets put on top of a tablet equipped with a pen that could both write in ink and sample coordinates of the tip and pressure exerted. The tablet was connected to a PC that recorded movements and displayed them in real-time. A software was used to extract 18 features for each task. Features and tasks are described in details in Appendix A. The dataset includes also a feature for identification of participants and another one that indicates whether the sample is associated to an AD patient (P) or an healthy person (H). This feature is used as label in the classification problem via a (classical or quantum) supervised learning model.
2.2 Support vector classification (SVC)
2.2.1 Classical SVC
Support vector machines (SVMs) are indeed supervised learning models for classification and regression [14]. Considering a binary classification problem in which the aim is to decide whether a set of samples belongs to one of two classes, each data point is viewed as a $d$-dimensional vector. The goal is to separate the points with a $(d-1)$-dimensional hyperplane such that the distance from the hyperplane to the nearest data points on each side, called support vectors, is maximized. A graphic representation of this is given in Fig. 2. In the original space this problem has often no solution, i.e. the sets of point are not linearly separable. A SVM uses the kernel trick, the technique of mapping inputs into high-dimensional feature spaces, to solve this problem [15]. The dot products of input pairs in the feature space are called kernels. Defining a map $\varphi:\mathcal{X}\rightarrow\mathcal{F}$ from the original input space to the feature space is equivalent to defining a kernel function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{C}$ that maps couples of inputs into their dot products in the feature space.
Figure 2: Graphic representation of a classification model in dimension $d=2$. The hyperplane is a straight line. Support vectors are circled.
For our classification task we consider 3 different kernel functions [16]. The radial basis function (RBF) kernel maps a pair of inputs $(x,x^{\prime})$ into $k_{rbf}(x,x^{\prime})=\exp\left(-\gamma|x-x^{\prime}|^{2}\right).$ This is one of the most common choices amongst kernel functions and it is used by default in the standard version of SVC implemented in the Scikit library in Python [17]. The linear kernel function is a simple dot product, $k_{lin}(x,x^{\prime})=\langle x,x^{\prime}\rangle.$ Lastly, the sigmoid kernel is given by $k_{sig}(x,x^{\prime})=\tanh(\gamma\langle x,x^{\prime}\rangle+r)$.
2.2.2 Quantum SVC
In the quantum version of SVMs, inputs are mapped into the space of quantum states [18]. The feature map is defined by a parametrized quantum circuit (PQC), whose parameters are given by the features of the inputs. A PQC is a sequence of parametrized unitary operators, called quantum logical gates, acting on a set of qubits, or quantum bits, which are two-state quantum systems [19]. To introduce the representation of qubit states, we need to define Dirac’s notation for unit vectors in $\mathbb{C}^{2}$. The standard basis of $\mathbb{C}^{2}$ is given by the vectors $\{|0\rangle,|1\rangle\},$ where $|0\rangle=\begin{pmatrix}1\\ 0\end{pmatrix}$, and $|1\rangle=\begin{pmatrix}0\\ 1\end{pmatrix}$. A generic unit vector in $\mathbb{C}^{2}$ is $|\phi\rangle=\alpha|0\rangle+\beta|1\rangle$, where $\alpha,\beta\in\mathbb{C}$ are such that $|\alpha|^{2}+|\beta|^{2}=1.$ The inner product of two unit vectors $|\phi\rangle,|\psi\rangle\in\mathbb{C}^{2}$ is denoted by $\langle\phi|\psi\rangle,$ while their outer product is $|\phi\rangle\langle\psi|.$ The state of a qubit is defined by a hermitian, positive semi-definite matrix $\rho\in\mathbb{C}^{2\times 2}$, such that its trace is equal to 1. The state is pure when there exists a $|\phi\rangle\in\mathbb{C}^{2}$ such that $\rho=|\phi\rangle\langle\phi|$, otherwise it is mixed.
To compute kernels we measure the fidelity of pairs of quantum states. Fidelity evaluates the degree of similarity of a pair of quantum states [20]. Given a pair of pure quantum states $\rho=|\phi\rangle\langle\phi|$ and $\sigma=|\psi\rangle\langle\psi|,$ their fidelity is defined by $\mathcal{F}(\rho,\sigma)=|\langle\phi|\psi\rangle|^{2}$. The most used generalization of this definition for a generic pair of mixed states $\rho$ and $\sigma$ is $\mathcal{F}(\rho,\sigma)=\left(\mathrm{tr}\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}}\right)^{2}$. The calculation of fidelities is the only computation requiring a quantum hardware. All other computations in quantum SVCs are handled by a classical hardware.
Quantum SVC is implemented using the standard version in Scikit library with precomputed kernels. We use different PQCs, scaling the number of qubits with similar ansatze. The ansatze are chosen considering the following aspects. The number of quantum gates needs to be small enough to make the time needed to perform the calculations as small as possible. Time of execution is a critical point in quantum computation, since it is often limited on real QPUs. Given a kernel function $k$, the Gram matrix associated to it is the matrix whose entries are $G_{ij}=k(x_{i},x_{j}),$ where $x_{i}$ and $x_{j}$ are two inputs. The eigenvalue curve of the matrix $G$ is flat when the problem is difficult to learn, whereas a desirable situation is found when the eigenvalue curve is non-flat and decays, either polynomially or exponentially fast [21]. An ansatz is expressive when it explores the space of unitaries as fully and uniformally as possible. Expressive ansatze are desirable when facing a problem with no prior knowledge of it, because of their ability to adapt to the task. However it is proved that expressive anstatze are difficult to train [22], while an ansatz with low expressibility is easier to train and can achieve a better performance on the task it is specialized to solve [23]. One possible solution to reduce expressibility is introducing a bandwidth factor, a scaling hyperparameter that multiplies the parameters in the PQC. The bandwidth parameter limits the reach of the feature map [24].
The PQCs we use are composed of $R_{x}$ and $R_{y}$ rotation gates, coontrolled-z and controlled-x gates. The gates $R_{x}$ and $R_{y}$ are 1-qubit parametrized unitaries that perform a rotation of an angle given by its parameter about the x-axis and the y-axis of the Bloch sphere, respectively. The Bloch sphere is a 3-dimensional geometrical representation of the states of a qubit. The matrix representations of the rotation gates are the following:
$R_{x}(\theta)=\begin{pmatrix}\cos\left(\frac{\theta}{2}\right)&-i\sin\left(\frac{\theta}{2}\right)\\ -i\sin\left(\frac{\theta}{2}\right)&\cos\left(\frac{\theta}{2}\right)\end{pmatrix},\quad R_{y}(\theta)=\begin{pmatrix}\cos\left(\frac{\theta}{2}\right)&-\sin\left(\frac{\theta}{2}\right)\\ \sin\left(\frac{\theta}{2}\right)&\cos\left(\frac{\theta}{2}\right)\end{pmatrix}.$
A controlled gate is a quantum operation acting on multiple qubits. One or more qubits act as controls, i.e. based on their configuration, a certain gate is applied or not to the remaining qubit, called target. The controlled-x gate is a 2-qubits controlled gate. When the control is in the state $|1\rangle$ an $X$ gate is applied to the target, where $X=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}.$ The controlled-z is also a 2-qubits controlled gate. It is symmetric, meaning that both qubits act as control, activated by the state $|1\rangle$, and target. The matrix representations of controlled-x and controlled-z gates are
$CX=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix},\quad CZ=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&-1\end{pmatrix}.$
A control qubit activated by the state $|1\rangle$ is represented in a circuit as $\bullet$.∙ . A target qubit to which an $X$ gate is applied is denoted by $\oplus$. Therefore the graphic representations of controlled-x and controlled-z in the circuits are, respectively, the following.
{quantikz}
[…]
In Fig. 3 we show the graphic representation of our 6-qubits circuit, composed by layers of $R_{x}$ and $R_{y}$ gates alternating with layers of controlled-z and controlled-x gates. On the right-hand side are symbols representing measurements applied to the qubits, converting the quantum states in a sequence of classical bits. The number of circuit parameters is 24.
[…]
Figure 3: Graphic representation of the 6-qubits PQC that maps a sample $s=(s[0],s[1],...,s[23])$ into a quantum state. The parameters in the rotation gates are given by a bandwidth factor $b=0.4$, chosen in order to maximize mean accuracy, times the components of $s$.
A similar structure with the same number of parameters is used to define also an 8-qubits and a 12-qubits circuit. The number of parameters must be equal to the number of features of the samples, so that each input feature corresponds to a parameter. Therefore the dataset needs to be processed through a principal component analysis (PCA) prior to applying the methods [25]. In order to reduce expressibility, the parameters in the rotation gates are multiplied by a bandwidth hyperparameter, chosen so as to maximize the accuracy reached by the classification. These circuits produce kernel matrices whose eigenvalue curve are plotted in Fig. 4. The curves are not flat but decay, hence the problem can be solved in this way.
Figure 4: Plot (in logarithmic scale) of the eigenvalue curves associated with the kernel matrices defined by the 6-qubits, 8-qubits and 12-qubits circuit described in the main text.
2.3 Implementation
Our algorithms are implemented in Python, using the functionalities of Scikit and Qiskit libraries [26]. First we perform some preprocessing on the dataset. In particular we drop the identification feature, while keeping the feature healthy/AD’s patient (H/P) as a label. We standardize the samples, by subtracting the mean and scaling to unit variance. Then we perform a PCA with 24 components. Finally we standardize the set of samples derived from the PCA. In fact, from our analysis it emerges that this second normalization improves the performance of both classical and quantum methods. To reduce the bias introduced by selecting the samples for training and test, we select randomly 20 different splittings of the dataset in training and test set, keeping for the latter 21% of the samples, so that 137 samples are used for training and 37 for test. We use the same splittings for all methods.
After that we run classical SVCs using 3 different kernel functions: a radial basis function, a linear function and a sigmoid function. Subsequently we run the quantum SVCs for each of the 3 PQCs. Besides imposing the choice of the kernel function in the classical SVCs and the option of precomputing the kernels in the quantum SVCs, we keep the default setting for the other options of command ‘SVC’. For each run we compute the accuracy score, i.e. the number of test samples labelled correctly divided by the total number of test samples, and finally we calculate and compare the obtained average accuracies.
Table 1: Accuracy (in percentage) of classical SVCs and quantum SVCs for 20 different splittings of the dataset in training and test set. The highest accuracy in each run is reported in red.
Classical SVC Quantum SVC
Rbf Linear Sigmoid 6 qubits 8 qubits 12 qubits
81.08 86.49 86.49 86.49 83.78 89.19
83.78 81.08 86.49 83.78 81.08 81.08
75.68 67.57 83.78 78.38 75.68 75.68
94.59 89.19 81.08 91.89 91.89 91.89
83.78 81.08 81.08 81.08 86.49 83.78
89.19 83.78 91.89 94.59 89.19 91.89
91.89 78.38 89.19 91.89 94.59 91.89
91.89 83.78 89.19 89.19 91.89 89.19
89.19 89.19 86.49 83.78 91.89 89.19
94.59 97.30 91.89 89.19 94.59 94.59
86.49 86.49 86.49 89.19 83.78 86.49
94.59 91.89 86.49 91.89 94.59 91.89
86.49 83.78 78.38 81.08 86.49 86.49
78.38 83.78 83.78 81.08 86.49 89.19
86.49 75.68 75.68 81.08 81.08 81.08
86.49 89.19 89.19 89.19 94.59 89.19
86.49 86.49 81.08 81.08 89.19 86.49
83.78 86.49 81.08 86.49 91.89 86.49
86.49 83.78 83.78 83.78 89.19 86.49
83.78 83.78 89.19 83.78 91.89 91.89
3 Results and Discussion
In Table 1 and in Fig. 5 we show the accuracy reached by the methods for each of the 20 splittings of the dataset in training and test set. Amongst classical SVCs, the one defined by the RBF kernel achieves average accuracy of 86.76%. Linear kernel and sigmoid kernel obtain 84.46% and 85.14% respectively. The quantum SVC defined by the 6-qubits PQC gets to 85.95%, the one defined by the 8-qubits PQC reaches 88.51% and the one defined by the 12-qubits circuit achieves average accuracy of 87.70%. One of the classical methods performs better than any other quantum method for 5 splittings. In 3 other splittings the same best precision is achieved by a classical and a quantum method. In the other 12 cases the quantum methods perform better. In particular, the 6-qubits method performs best in 2 splittings, the 8-qubits method in 7 splittings, the 12-qubits method achieves the best precision in 2 cases, while it obtains the same best accuracy of the 8-qubits method in 1 other case.
Figure 5: Accuracies (in percentage) vs. splittings of the dataset in classical and quantum SVC models.
3.1 Data subsampling
Now we repeat the analysis above when restricted to only one category of features, e.g. the features derived only from graphic tasks or only copy tasks or only memory and dictation tasks. We compare the performance of the classical and quantum SVCs that have achieved the highest average accuracies above, i.e. the classical SVC defined by the RBF kernel and the 8-qubits quantum SVC.
Table 2: Accuracies (in percentage) of RBF classical SVC and 8-qubits SVC achieved combining with the majority vote decision rule the predictions made considering all tasks individually (left) and considering only the best 5 tasks (right). The highest accuracy in each run is reported in red.
All tasks Best 5 tasks
Classical Quantum Classical Quantum
78.38 81.08 89.19 91.89
81.08 78.38 86.49 86.49
72.97 72.97 86.49 83.78
78.38 81.08 91.89 91.89
81.08 81.08 91.89 91.89
78.38 81.08 81.08 83.78
89.19 89.19 86.49 86.49
72.97 75.68 81.08 83.78
83.78 86.49 91.89 91.89
91.89 94.59 94.59 94.59
89.19 83.78 89.19 89.19
70.27 72.97 89.19 89.19
75.68 72.97 81.08 83.78
78.38 78.38 81.08 81.08
81.08 81.08 78.38 78.38
78.38 83.78 83.78 83.78
89.19 91.89 89.19 86.49
86.49 83.78 94.59 89.19
75.68 78.38 89.19 89.19
75.68 81.08 86.49 86.49
We use the same 20 splittings of the dataset in training and test set we have used above. First we take into account the features derived from the tasks in the category ‘graphic’. RBF and 8-qubits SVC achieve average accuracies of 79.19% and 79.86% respectively. After that we consider only the features derived from the tasks in the category ‘copy’. Both classical and quantum SVC achieve average accuracies of 83.11%. Finally, we focus on the features derived from the tasks in the category ‘memory and dictation’. Classical and quantum SVC achieve mean accuracies of 79.46% and 78.38%.
Subsequently we consider the features obtained from each of the 25 task individually. For each task we run RBF SVC, store the predictions made by the 25 models and use the majority vote decision rule [27] to select the outputs; namely for each sample we select the output predicted by most of the models. Likewise we run for each task a quantum SVC. Since each task is described by 18 features, we need to define a new PQC with 18 parameters. For this purpose we choose a 9-qubits circuit. No PCA is necessary in this case, since the number of parameters coincides with the number of features. We perform 20 runs with the same splittings in training and test set we have used above. The accuracies reached at each run are shown in Table 2. Classical and quantum SVC obtain mean accuracies of 80.41% and 81.49% respectively. Afterwards we select the best 5 tasks, namely the tasks that have achieved singularly the best accuracies in the previous runs. For both classical and quantum SVC the best tasks, although in a different order, are 23, 19, 17, 24 and 8. The tasks are described in Appendix A. As before we use the majority vote decision rule to select the predictions amongst those made by the 5 models associated with the tasks. In this case both classical and quantum SVC obtain average accuracies of 87.16%.
3.2 Noise robustness
Due to limited accessibility to real quantum hardware, computation of kernels is simulated without noise on a classical hardware, using tools from Qiskit library. To show how noise affects our results, we implement some models introducing depolarizing, bit-flip and amplitude damping errors [28].
First we consider the quantum SVC defined by the 6-qubits PQC. We select a splitting of the dataset in training and test set and we run the method without any noise. Then we add noise through the depolarizing error model, initially with error rate $p=0.01$, and after that we progressively increase $p$ by 0.01. For each run we compute the achieved accuracy. We repeat it also for bit-flip error and amplitude damping error, and subsequently we perform the same analysis for the 8-qubits and the 12-qubits methods. For the latter methods, excessive computational cost makes it necessary to reduce the number of samples before running the methods. In particular we consider 20 samples per label for the 8-qubits method, and 15 samples per label for the 12-qubits method.
In Fig. 6 we show how the accuracies of the 6-qubits SVC vary as the noise rate $p$ increases in the 3 different noise models. The accuracy curves stop at the $p$ value where the methods collapse, namely at $p^{*}$ such that for all $p\in[p^{*},1]$ the accuracy is the same as the one obtained at $p^{*}$. In the 8-qubits and 12-qubits SVCs, since the amount of noise injected grows with the number of qubits, the collapse happens for smaller $p^{*}$. All methods seems to be more robust to amplitude damping error than to depolarizing and bit-flip errors.
Figure 6: Accuracies (in percentage) of 6-qubits SVC as noise rate increases in the depolarizing, bit-flip and amplitude damping noise models.
4 Conclusion
The aim of this work is comparing the performances of classical and quantum kernel methods for classifying samples of AD patients and healthy people subjected to handwriting tests. To achieve this, we have used first some classical approaches and then we have compared their performances with the ones achieved by quantum methods defined by various PQCs. Amongst classical SVCs, the one defined by the radial basis function kernel achieves average accuracy of 86.76%, while linear kernel and sigmoid kernel obtain 84.46% and 85.14% respectively. The quantum SVC defined by the 6-qubits PQC gets to 85.95%, the one defined by the 8-qubits PQC reaches 88.51% and the one defined by the 12-qubits circuit achieves average accuracy of 87.70%. The fact that the latter does not outperform the model defined by the 8-qubits circuit seems to indicate that performance does not automatically improve by increasing the number of qubits, and some ansatze are better suited for smaller PQCs, hence more feasible with the currently available quantum processors.
In addition to running the methods considering all features together we have also defined methods based on data subsampling. We have considered separately the features associated with the 3 categories of tasks, namely features derived from graphic tasks, copy tasks or memory and dictation tasks. With graphic features, RBF and 8-qubits SVC achieve average accuracies of 79.19% and 79.86% respectively. When considering copy features, both RBF and 8-qubits SVC achieve average accuracy of 83.11%. Using the features derived from the tasks in the category ‘memory and dictation’, RBF and 8-qubits SVC achieve average accuracies of 73.65% and 71.89% respectively. Then we have examined the features obtained from each of the 25 task individually, defining a model for each task, and we have used the majority vote decision rule to combine the predictions made by the 25 models. Classical and quantum SVC obtain mean accuracies of 80.41% and 81.49% respectively. When restricted to only the best 5 tasks, i.e. the tasks that have achieved singularly the best accuracies in the previous runs, both classical and quantum SVC obtain average accuracies of 87.16%.
Kernels in quantum methods have been calculated simulating quantum computation without noise. We have also investigated how different types of noise affect the performance of quantum methods. From our analysis it appears that the methods show good robustness to the error introduced by amplitude damping, which is associated to energy dissipation. In fact, quantum SVCs tend to perform poorly only for high value of the noise rate related to the damping probability. Both depolarization and bit flip error instead cause the methods to perform worse for smaller values of noise rate. However, on real quantum computers different sources of noise act simultaneously, due to interaction with the environment and interaction between qubits, or imperfect application of quantum gates [29]. Therefore quantum noise models, without in-depth physical experiments, can only partially tackle the complexity of this problem.
Our study shows that quantum methods can achieve better performances in classification tasks even involving big datasets, as for the case of medical data analysis. Nowadays only QPUs with a limited number of noisy qubits are available. However this does not seem to affect too much the capability of quantum kernel methods. Moreover the constant improvement of quantum technologies suggests that there is a lot of room for enhancements to make quantum methods achieve even better results.
The techniques used in our work may have a wide range of applications in QML algorithms for classification and regression, and can be implemented to enhance classical ML models for similar problems of early screening, e.g. detection of autism in children [30], or identification of various types of cancer [31, 32].
Data availability
The DARWIN dataset is available at https://archive.ics.uci.edu/dataset/732/darwin.
Code availability
The code used is available at request.
Acknowledgements
This work was supported by the European Commission’s Horizon Europe Framework Programme under the Research and Innovation Action GA n. 101070546–MUQUABIS, by the European Union’s Horizon 2020 research and innovation programme under FET-OPEN GA n. 828946–PATHOS, by the European Defence Agency under the project Q-LAMPS Contract No B PRJ- RT-989, and by the MUR Progetti di Ricerca di Rilevante Interesse Nazionale (PRIN) Bando 2022 - project n. 20227HSE83 – ThAI-MIA funded by the European Union - Next Generation EU.
G. C. is a member of INFM (INdAM).
Appendix A Dataset details
The 25 handwriting/drawing tasks can be grouped in 3 categories: memory and dictation, graphic, copy. The tasks are described in Table 3.
Table 3: List of tasks performed. The tasks are divided in the categories memory and dictation (M), graphic (G), and copy (C).
# Description Category
1 Signature drawing M
2 Join two points with a horizontal line, continuously for four times G
3 Join two points with a vertical line, continuously for four times G
4 Retrace a circle (6 cm of diameter) continuously for four times G
5 Retrace a circle (3 cm of diameter) continuously for four times G
6 Copy the letters ‘l’, ‘m’ and ‘p’ C
7 Copy the letters on the adjacent rows C
8 Write cursively a sequence of four lowercase letter ‘l’, in a single smooth movement C
9 Write cursively a sequence of four lowercase cursive bigram ‘le’, in a single smooth movement C
10 Copy the word “sheet” C
11 Copy the word “sheet” above a line C
12 Copy the word “mum” C
13 Copy the word “mum” above a line C
14 Memorize the words “telephone”, “dog”, and “shop” and rewrite them M
15 Copy in reverse the word “bottle” C
16 Copy in reverse the word “house” C
17 Copy six words (regular, non regular, non words) in the appropriate boxes C
18 Write the name of the object shown in a picture (a chair) M
19 Copy the fields of a postal order C
20 Write a simple sentence under dictation M
21 Retrace a complex form G
22 Copy a telephone number C
23 Write a telephone number under dictation M
24 Draw a clock, with all hours and put hands at 11:05 (Clock Drawing Test) G
25 Copy a paragraph C
[…]
Each task is performed on a different paper sheet and the pile of 25 sheets is put on top of a Wacom’s Bamboo tablet equipped with a pen that can both write in ink on paper and allow the tablet to sample x-y coordinate on paper and on air within a maximum distance of 3cm. The tablet is connected to a PC that processes the raw data to obtain for each task the following features:
  1. 1.
    Total Time (TT): Total time spent to perform the entire task.
  2. 2.
    Air Time (AT): Time spent to perform in-air movements.
  3. 3.
    Paper Time (PT): Time spent to perform on-paper movements.
  4. 4.
    Mean Speed on-paper (MSP): Average speed of on-paper movements.
  5. 5.
    Mean Speed in-air (MSA): Average speed of in-air movements.
  6. 6.
    Mean Acceleration on-paper (MAP): Average acceleration of on-paper movements.
  7. 7.
    Mean Acceleration in-air (MAA): Average acceleration of in-air movements.
  8. 8.
    Mean Jerk on-paper (MJP): Average jerk of on-paper movements. Jerk is the variation of acceleration with respect to time.
  9. 9.
    Mean Jerk in-air (MJA): Average jerk of in-air movements.
  10. 10.
    Pressure Mean (PM): Average of the pressure levels exerted by the pen tip.
  11. 11.
    Pressure Var (PV): Variance of the pressure levels exerted by the pen tip.
  12. 12.
    GMRT on-paper (GMRTP): Generalization of the Mean Relative Tremor (MRT) as defined by Pereira et al. (2015). MRT measures the amount of tremor in drawing spirals and meanders.
  13. 13.
    GMRT in-air (GMRTA): Generalization of the Mean Relative Tremor computed on in air movements.
  14. 14.
    Mean GMRT (GMRT): Average of GMRTP and GMRTA.
  15. 15.
    Pendowns Number (PWN): Counts the total number of pendowns recorded during the execution of the entire task.
  16. 16.
    Max X Extension (XE): Maximum extension recorded along the X axis.
  17. 17.
    Max Y Extension (YE): Maximum extension recorded along the Y axis.
  18. 18.
    Dispersion Index (DI): The Dispersion Index measures how the handwritten trace is “dispersed” on the entire piece of paper. To calculate the index the sheet is ideally divided into TB boxes of 3 × 3 pixels, then the number CB of boxes containing a fragment of handwriting/drawing is computed. DI is given by the ratio between CB and TB.

### v2

@@
 Quantum machine learning is a new research field combining quantum information science and machine learning.
-Quantum computing technologies seem to be particularly well suited to solving problems in the health sector in an efficient way, because they may deal with large datasets more efficiently than classical AI.
+Quantum computing technologies appear to be particularly well-suited for addressing problems in the health sector efficiently.
+They have the potential to handle large datasets more effectively than classical models and offer greater transparency and interpretability for clinicians.
 Alzheimer’s disease is a neurodegenerative brain disorder that mostly affects elderly people, causing important cognitive impairments.
@@
 Here we apply quantum AI to this use-case.
-In particular, we use this dataset to test kernel methods for classification task and compare their performances with the ones obtained via quantum machine learning methods.
-We find that quantum and classical algorithms achieve similar performances and in some cases quantum methods perform even better.
+In particular, we use this dataset to test classical methods for classification and compare their performances with the ones obtained via quantum machine learning methods.
+We find that quantum methods generally perform better than classical methods.
 Our results pave the way for future new quantum machine learning applications in early-screening diagnostics in the healthcare domain.
@@
 Since there is no cure for AD, it is critical to improve the approaches now used for diagnosis [3].
+Early diagnosis can make patients eligible for clinical trials, which are essential for the development of new treatments.
+Participation in these trials can provide access to cutting-edge therapies that are not yet widely available [4] and it can help maintain cognitive function for a longer period, allowing individuals to live more independently and enjoy a better quality of life [5].
+Many techniques have been used for this purpose.
+Neuroimaging techniques are commonly used to detect brain atrophy and other structural changes associated with AD; however their limitations include restricted availability and the inability to detect early-stage diseases [6].
+Cognitive assessments, on the other hand, despite potentially being influenced by the patient’s education level, measure abilities such as memory, attention, language, and problem-solving, while motor tests examine coordination, balance, and gait.
+This comprehensive approach aids in fully understanding the extent of the disease’s impact [7].
 Handwriting is a task that involves both cognitive and motor functions to plan and properly execute the movements that are required with high coordination [8].
@@
 [10] it is described how this dataset can be exploited by machine learning models to discriminate between AD patients and healthy people.
-Machine learning (ML) is the field of artificial intelligence (AI) that studies the development of learning models to recognize, for example, patterns in large sets of data [7].
-These models can learn how to attribute outputs to unseen inputs for classification or regression tasks (supervised learning), or find a structure in the inputs (unsupervised learning) for clustering or dimensionality reduction [8].
-A lot of work has been done in the past years to improve the performance of ML algorithms and make them capable of handling efficiently big datasets.
-One more recent and promising solution could be to use quantum machine learning (QML) techniques.
-In QML inputs are encoded in quantum states and computation is done on a quantum computer [9, 10].
+Machine learning (ML) is a subfield of artificial intelligence (AI) that focuses on the development of algorithms and statistical models that enable computers to learn and make predictions or decisions based on data [11].
+These models can identify patterns, trends, and relationships within large datasets, allowing for tasks such as classification, regression, clustering, and anomaly detection [12].
+One of the main obstacles to the usage of AI in the healthcare field, and in particular for the diagnosis of diseases, is the “black box” nature of many AI models.
+Medical practitioners often find it challenging to trust AI predictions because they cannot easily understand how these models arrive at their conclusions [13].
+Quantum AI can improve the transparency and interpretability of AI models by leveraging the principles of quantum mechanics and providing a clear visual representation of models [14], allowing clinicians to understand the decision-making process better.
+Additionally, it can handle large volumes of data more efficiently than classical AI, making it a powerful tool for data-intensive tasks [15].
+In quantum machine learning (QML) inputs are encoded in quantum states and computation is done on a quantum computer [16, 17].
 Although quantum processing units (QPUs) are still subjected to relatively large noise rates, it seems that applications in ML can lead to remarkable results.
@@
 [19].
-A previous study [13] approaches via a so-called variational quantum classifier (VQC) the problem of classification defined by the same dataset we consider in our work.
+A previous study [20] approaches via a variational quantum classifier (VQC) the problem of classification defined by the same dataset we consider in our work.
 Here we focus in particular on quantum kernel methods, i.e.
 kernel methods in which the feature space is a space of quantum states.
-We use these algorithms to classify the samples and compare their performances with the ones achieved by classical algorithms.
+We utilize these algorithms to classify the samples and evaluate their performance against a classical kernel algorithm and two additional classical methods.
 The paper is organized as follows.
-In Section 2 we describe how the DARWIN dataset was created and we define the classical and quantum kernel methods for classification.
+In Section 2 we describe how the DARWIN dataset was created and we define the classical and quantum methods for classification.
 Then in Section 3 and 4 we report their performances, we discuss the results and draw our conclusions.
@@
 [9].
-It includes 25 tasks that can be divided in the following categories: graphic tasks, like joining points and drawing geometrical figures; copy tasks, testing the ability of repeating more complex gestures to write letters, numbers and words; memory tasks, involving writing previously memorized words; and dictation tasks.
+This protocol includes 25 tasks that can be divided in the following categories: graphic tasks, like joining points and drawing geometrical figures; copy tasks, testing the ability of repeating more complex gestures to write letters, numbers and words; memory tasks, involving writing previously memorized words; and dictation tasks.
 The protocol was submitted to 174 participants: 89 AD patients and 85 healthy people, recruited so that the two groups would match in terms of age, education and gender.
@@
 This feature is used as label in the classification problem via a (classical or quantum) supervised learning model.
-2.2 Support vector classification (SVC)
-2.2.1 Classical SVC
-Support vector machines (SVMs) are indeed supervised learning models for classification and regression [14].
+2.2 Classical SVC
+Support vector machines (SVMs) are supervised learning models for classification and regression [21].
 Considering a binary classification problem in which the aim is to decide whether a set of samples belongs to one of two classes, each data point is viewed as a $d$-dimensional vector.
@@
 Figure 2: Graphic representation of a classification model in dimension $d=2$. The hyperplane is a straight line. Support vectors are circled.
-For our classification task we consider 3 different kernel functions [16].
+For our classification task we consider different kernel functions [23].
 The radial basis function (RBF) kernel maps a pair of inputs $(x,x^{\prime})$ into $k_{rbf}(x,x^{\prime})=\exp\left(-\gamma|x-x^{\prime}|^{2}\right).$ This is one of the most common choices amongst kernel functions and it is used by default in the standard version of SVC implemented in the Scikit library in Python [24].
-The linear kernel function is a simple dot product, $k_{lin}(x,x^{\prime})=\langle x,x^{\prime}\rangle.$ Lastly, the sigmoid kernel is given by $k_{sig}(x,x^{\prime})=\tanh(\gamma\langle x,x^{\prime}\rangle+r)$.
-2.2.2 Quantum SVC
+The linear kernel function is a simple dot product, $k_{lin}(x,x^{\prime})=\langle x,x^{\prime}\rangle.$ The polynomial function is $k_{pol}(x,x^{\prime})=(\gamma\langle x,x^{\prime}\rangle+r)^{d}.$ Lastly, the sigmoid kernel is given by $k_{sig}(x,x^{\prime})=\tanh(\gamma\langle x,x^{\prime}\rangle+r)$.
+2.3 k-Nearest Neighbors
+The k-Nearest Neighbors (kNN) model is a method used both for classification and regression [25].
+With this technique, a label is assigned to a data point based on the most common label among its $k$ nearest neighbors.
+In particular, the probability that the unknown sample $x_{0}$ belongs to class $j$ is given by
+$P(Y=j\mid X=x_{0})=\frac{1}{k}\sum_{i\in N_{0}}I(y_{i}=j),$
+where $k$ is the number of nearest neighbors, $N_{0}$ is the set of the $k$ nearest neighbors to $x_{0}$ and $I(y_{i}=j)$ is an indicator function that equals 1 if the $i$-th neighbor $y_{i}$ has the label $j$, and 0 otherwise.
+2.4 Decision Tree
+A Decision Tree (DT) is a decision support tool that visually represents decisions and their possible outcomes in a hierarchical, tree-like structure [26]. This structure is composed of:
+  • nodes, which are decision points where a specific feature is tested;
+  • branches, representing the possible outcomes of the test conducted at each node;
+  • leaf nodes where a final decision is made and a label is assigned.
+The tree is constructed by recursively splitting the data based on the attribute that provides the best separation according to a chosen criterion (e.g., Gini impurity or information gain).
+This process continues until the stopping criteria are met, such as a maximum tree depth or a minimum number of samples per leaf node.
+2.5 Quantum SVC
 In the quantum version of SVMs, inputs are mapped into the space of quantum states [27].
@@
 Therefore the graphic representations of controlled-x and controlled-z in the circuits are, respectively, the following.
-{quantikz}
-[…]
-[…]
 In Fig.
-3 we show the graphic representation of our 6-qubits circuit, composed by layers of $R_{x}$ and $R_{y}$ gates alternating with layers of controlled-z and controlled-x gates.
+3 we show the graphic representation of a 6-qubits circuit, composed by layers of $R_{x}$ and $R_{y}$ gates alternating with layers of controlled-z and controlled-x gates.
 On the right-hand side are symbols representing measurements applied to the qubits, converting the quantum states in a sequence of classical bits.
 The number of circuit parameters is 24.
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Figure 3: Graphic representation of the 6-qubits PQC that maps a sample $s=(s[0],s[1],...,s[23])$ into a quantum state.
-The parameters in the rotation gates are given by a bandwidth factor $b=0.4$, chosen in order to maximize mean accuracy, times the components of $s$.
+Figure 3: Graphic representation of a 6-qubits PQC that maps a sample $s=(s[0],s[1],...,s[23])$ into a quantum state.
+The parameters in the rotation gates are given by a bandwidth factor $b=0.4$, chosen in order to maximize the accuracy, times the components of $s$.
 A similar structure with the same number of parameters is used to define also an 8-qubits and a 12-qubits circuit.
@@
 Therefore the dataset needs to be processed through a principal component analysis (PCA) prior to applying the methods [34].
-In order to reduce expressibility, the parameters in the rotation gates are multiplied by a bandwidth hyperparameter, chosen so as to maximize the accuracy reached by the classification.
+In order to reduce expressibility, the parameters in the rotation gates are multiplied by a tunable bandwidth hyperparameter.
 These circuits produce kernel matrices whose eigenvalue curve are plotted in Fig.
 4.
-The curves are not flat but decay, hence the problem can be solved in this way.
+Since the curves decay, the problem can be addressed using this approach.
 Figure 4: Plot (in logarithmic scale) of the eigenvalue curves associated with the kernel matrices defined by the 6-qubits, 8-qubits and 12-qubits circuit described in the main text.
@@
 Finally we standardize the set of samples derived from the PCA.
-In fact, from our analysis it emerges that this second normalization improves the performance of both classical and quantum methods.
-To reduce the bias introduced by selecting the samples for training and test, we select randomly 20 different splittings of the dataset in training and test set, keeping for the latter 21% of the samples, so that 137 samples are used for training and 37 for test.
-We use the same splittings for all methods.
-After that we run classical SVCs using 3 different kernel functions: a radial basis function, a linear function and a sigmoid function.
-Subsequently we run the quantum SVCs for each of the 3 PQCs.
-Besides imposing the choice of the kernel function in the classical SVCs and the option of precomputing the kernels in the quantum SVCs, we keep the default setting for the other options of command ‘SVC’.
-For each run we compute the accuracy score, i.e.
-the number of test samples labelled correctly divided by the total number of test samples, and finally we calculate and compare the obtained average accuracies.
-Table 1: Accuracy (in percentage) of classical SVCs and quantum SVCs for 20 different splittings of the dataset in training and test set. The highest accuracy in each run is reported in red.
-Classical SVC Quantum SVC
-Rbf Linear Sigmoid 6 qubits 8 qubits 12 qubits
-81.08 86.49 86.49 86.49 83.78 89.19
-83.78 81.08 86.49 83.78 81.08 81.08
-75.68 67.57 83.78 78.38 75.68 75.68
-94.59 89.19 81.08 91.89 91.89 91.89
-83.78 81.08 81.08 81.08 86.49 83.78
-89.19 83.78 91.89 94.59 89.19 91.89
-91.89 78.38 89.19 91.89 94.59 91.89
-91.89 83.78 89.19 89.19 91.89 89.19
-89.19 89.19 86.49 83.78 91.89 89.19
-94.59 97.30 91.89 89.19 94.59 94.59
-86.49 86.49 86.49 89.19 83.78 86.49
-94.59 91.89 86.49 91.89 94.59 91.89
-86.49 83.78 78.38 81.08 86.49 86.49
-78.38 83.78 83.78 81.08 86.49 89.19
-86.49 75.68 75.68 81.08 81.08 81.08
-86.49 89.19 89.19 89.19 94.59 89.19
-86.49 86.49 81.08 81.08 89.19 86.49
-83.78 86.49 81.08 86.49 91.89 86.49
-86.49 83.78 83.78 83.78 89.19 86.49
-83.78 83.78 89.19 83.78 91.89 91.89
+In fact, from our analysis it emerges that normalizing twice improves the performance of both classical and quantum methods.
+The dataset is divided for cross-validation using ShuffleSplit, performing 20 splits: 60% of the data is used for training, 20% for validation and 20% as test set.
+For each method, a parameter grid is defined for hyperparameter tuning: in this phase the model is trained on the training set and evaluated on the validation set.
+After that, the model with the best hyperparameters is trained on the union of training and validation set and is evaluated on the test set.
+The mean and standard deviation of the accuracies across the splits are calculated.
+For classical SVC, the parameter grid includes different options for 4 hyperparameters:
+  • ‘kernel’, that defines the kernel function;
+  • ‘C’, used to balance error minimization and model complexity;
+  • ‘gamma’, which determines the influence range of training examples;
+  • ‘tol’, which sets the precision for stopping criteria.
+Two more classical methods for classifications are used for comparison, kNN and DT.
+For kNN the parameter grid includes 3 hyperparametrs: number of neighbors, metric and type of weights.
+For DT it includes 5 hyperparameters: separation criterion, splitter, maximum depth, minimum samples for split and minimum samples for leaf.
+For quantum SVCs, we use fixed ansatze with 6, 8, and 12 qubits.
+To simplify the computations, the only hyperparameter tuned during the validation phase is the bandwidth parameter.
+Therefore, except for selecting the option of running the model with precomputed kernels, the other parameters of SVC() are left by default: $\mathrm{C}=1.0$, ‘gamma’=‘scale’, i.e.
+it scales with the number of samples, and ‘tol’=0.001.
+The quantum methods are simulated locally using Aer Simulator.
 3 Results and Discussion
-In Table 1 and in Fig.
-5 we show the accuracy reached by the methods for each of the 20 splittings of the dataset in training and test set.
-Amongst classical SVCs, the one defined by the RBF kernel achieves average accuracy of 86.76%.
-Linear kernel and sigmoid kernel obtain 84.46% and 85.14% respectively.
-The quantum SVC defined by the 6-qubits PQC gets to 85.95%, the one defined by the 8-qubits PQC reaches 88.51% and the one defined by the 12-qubits circuit achieves average accuracy of 87.70%.
-One of the classical methods performs better than any other quantum method for 5 splittings.
-In 3 other splittings the same best precision is achieved by a classical and a quantum method.
-In the other 12 cases the quantum methods perform better.
-In particular, the 6-qubits method performs best in 2 splittings, the 8-qubits method in 7 splittings, the 12-qubits method achieves the best precision in 2 cases, while it obtains the same best accuracy of the 8-qubits method in 1 other case.
-Figure 5: Accuracies (in percentage) vs. splittings of the dataset in classical and quantum SVC models.
+In Table 1 we show the mean accuracy and standard deviation of classical and quantum methods.
+Amongst classical methods, SVC has the highest mean accuracy, but the standard deviation is relatively high, suggesting that its performance can vary significantly across different splits; kNN and DT are less effective in terms of average performance, although the latter offers more consistent results.
+The 6-qubits and 8-qubits SVCs perform comparably to the classical SVC, with slightly lower accuracy but a smaller standard deviation, indicating more consistent performance.
+The quantum SVC with 12 qubits outperforms all other models, both classical and quantum, with the highest accuracy and best consistency.
+This result suggests that increasing the number of qubits can enhance the performance of quantum classifiers.
+Table 1: Mean accuracy and standard deviation (expressed in percentage) of classical SVC, k-Nearest Neighbors, decision tree and quantum SVCs defined by 6-qubits, 8-qubits and 12-qubits PQCs.
+Classical Quantum
+SVC kNN DT 6q 8q 12q
+$85.28\pm 7.46$ $69.57\pm 8.20$ $73.57\pm 5.26$ $83.57\pm 4.59$ $83.14\pm 4.51$ $88.29\pm 4.69$
 3.1 Data subsampling
-Now we repeat the analysis above when restricted to only one category of features, e.g.
-the features derived only from graphic tasks or only copy tasks or only memory and dictation tasks.
-We compare the performance of the classical and quantum SVCs that have achieved the highest average accuracies above, i.e.
-the classical SVC defined by the RBF kernel and the 8-qubits quantum SVC.
-Table 2: Accuracies (in percentage) of RBF classical SVC and 8-qubits SVC achieved combining with the majority vote decision rule the predictions made considering all tasks individually (left) and considering only the best 5 tasks (right).
-The highest accuracy in each run is reported in red.
-All tasks Best 5 tasks
-Classical Quantum Classical Quantum
-78.38 81.08 89.19 91.89
-81.08 78.38 86.49 86.49
-72.97 72.97 86.49 83.78
-78.38 81.08 91.89 91.89
-81.08 81.08 91.89 91.89
-78.38 81.08 81.08 83.78
-89.19 89.19 86.49 86.49
-72.97 75.68 81.08 83.78
-83.78 86.49 91.89 91.89
-91.89 94.59 94.59 94.59
-89.19 83.78 89.19 89.19
-70.27 72.97 89.19 89.19
-75.68 72.97 81.08 83.78
-78.38 78.38 81.08 81.08
-81.08 81.08 78.38 78.38
-78.38 83.78 83.78 83.78
-89.19 91.89 89.19 86.49
-86.49 83.78 94.59 89.19
-75.68 78.38 89.19 89.19
-75.68 81.08 86.49 86.49
-We use the same 20 splittings of the dataset in training and test set we have used above.
-First we take into account the features derived from the tasks in the category ‘graphic’.
-RBF and 8-qubits SVC achieve average accuracies of 79.19% and 79.86% respectively.
-After that we consider only the features derived from the tasks in the category ‘copy’.
-Both classical and quantum SVC achieve average accuracies of 83.11%.
-Finally, we focus on the features derived from the tasks in the category ‘memory and dictation’.
-Classical and quantum SVC achieve mean accuracies of 79.46% and 78.38%.
-Subsequently we consider the features obtained from each of the 25 task individually.
-For each task we run RBF SVC, store the predictions made by the 25 models and use the majority vote decision rule [27] to select the outputs; namely for each sample we select the output predicted by most of the models.
-Likewise we run for each task a quantum SVC.
+Now we compare the performance of classical and quantum methods when restricted to subsets of features.
+For simplicity, we consider only the classical SVC and 12-qubits quantum SVC.
+To properly evaluate the performance of the methods on feature subsets, we use for both methods the hyperparameters that have been most frequently chosen in the previous analysis.
+First we use one category of features at a time, e.g.
+the features derived only from graphic tasks, then only copy tasks, and finally only memory and dictation tasks.
+The dataset is divided for cross-validation using ShuffleSplit, performing 20 splits: 80% of the data is used for training and 20% for testing.
+In Table 2 we show the mean accuracy and standard deviation of classical and quantum methods.
+Table 2: Mean accuracy and standard deviation (expressed in percentage) of classical SVC and 12-qubits SVC models applied to the subset of features obtained through copy, graphic or memory and dictation tasks.
+Classical 12q
+Copy $85.57\pm 8.00$ $85.71\pm 7.06$
+Graphic $78.14\pm 4.97$ $81.29\pm 5.68$
+Memory $79.28\pm 6.19$ $78.57\pm 4.99$
+Subsequently we consider the features obtained from each of the 25 tasks individually.
+For each task we run classical SVC, store the predictions made by the 25 models and use the majority vote decision rule [36] to select the outputs; namely for each sample we select the output predicted by most of the models.
 Since each task is described by 18 features, we need to define a new PQC with 18 parameters.
 For this purpose we choose a 9-qubits circuit.
-No PCA is necessary in this case, since the number of parameters coincides with the number of features.
-We perform 20 runs with the same splittings in training and test set we have used above.
-The accuracies reached at each run are shown in Table 2.
-Classical and quantum SVC obtain mean accuracies of 80.41% and 81.49% respectively.
-Afterwards we select the best 5 tasks, namely the tasks that have achieved singularly the best accuracies in the previous runs.
-For both classical and quantum SVC the best tasks, although in a different order, are 23, 19, 17, 24 and 8.
-The tasks are described in Appendix A.
+After considering all 25 tasks, we select the 5 tasks that achieve singularly the best accuracies in the previous runs.
+For classical SVC the tasks are 21, 17, 16, 7, 23, while for quantum SVC they are 21, 17, 24, 14, 23.
 As before we use the majority vote decision rule to select the predictions amongst those made by the 5 models associated with the tasks.
-In this case both classical and quantum SVC obtain average accuracies of 87.16%.
+In Table 3 we show the mean accuracy and standard deviation of classical and quantum methods over 20 splittings.
+Table 3: Mean accuracy and standard deviation (expressed in percentage) of classical SVC and 9-qubits SVC models achieved combining with the majority vote decision rule the predictions made considering all tasks individually and considering only the best 5 tasks.
+Classical 9q
+All 25 $85.71\pm 3.59$ $86.00\pm 4.30$
+Best 5 $80.28\pm 4.23$ $81.35\pm 5.19$
+Across all data subsampling scenarios, the quantum algorithm consistently demonstrates improved or comparable performance, underscoring its robustness and potential benefits in managing diverse data types.
 3.2 Noise robustness
-Due to limited accessibility to real quantum hardware, computation of kernels is simulated without noise on a classical hardware, using tools from Qiskit library.
-To show how noise affects our results, we implement some models introducing depolarizing, bit-flip and amplitude damping errors [28].
-First we consider the quantum SVC defined by the 6-qubits PQC.
-We select a splitting of the dataset in training and test set and we run the method without any noise.
-Then we add noise through the depolarizing error model, initially with error rate $p=0.01$, and after that we progressively increase $p$ by 0.01.
-For each run we compute the achieved accuracy.
-We repeat it also for bit-flip error and amplitude damping error, and subsequently we perform the same analysis for the 8-qubits and the 12-qubits methods.
-For the latter methods, excessive computational cost makes it necessary to reduce the number of samples before running the methods.
-In particular we consider 20 samples per label for the 8-qubits method, and 15 samples per label for the 12-qubits method.
+Due to limited accessibility to real quantum hardware, computation of kernels is simulated without noise on a classical hardware.
+Now we compare the performance of the methods executed on a noiseless Aer Simulator with those run on a fake backend.
+A fake backend is a simulated quantum computing environment that mimics the behavior of real IBM Quantum systems, providing real hardware qubit properties to simulate realistic conditions.
+We select a splitting of the dataset in training and test set, leaving 20% of the samples for the latter, and we run the quantum SVCs on the Aer Simulator without noise, with number of shots equal to 256.
+This means that each circuit run is executed 256 times and the measurement results are aggregated to provide a more accurate representation of the quantum state.
+The default number of shots is usually set to 1024, but we choose to reduce this to speed up the computation, especially when using the fake backend.
+Then, with the same splitting and the same number of shots, we run the methods on Fake Melbourne backend, which simulates a 14-qubits quantum processor.
+The average T1 and T2 times of Fake Melbourne are 50 and 70 microseconds respectively, where T1 is a measure of how long a qubit can retain its state before it decays to ground state, and T2 indicates how long a qubit can stay in a superposition state before losing coherence, due to interactions with the environment [37].
+A noiseless backend always yields the same result, whereas the performance of a fake backend, influenced by noise, is variable.
+Therefore we run the methods on the fake backend 20 times, storing their predictions and using the majority rule to yield the final prediction.
 In Fig.
-6 we show how the accuracies of the 6-qubits SVC vary as the noise rate $p$ increases in the 3 different noise models.
-The accuracy curves stop at the $p$ value where the methods collapse, namely at $p^{*}$ such that for all $p\in[p^{*},1]$ the accuracy is the same as the one obtained at $p^{*}$.
-In the 8-qubits and 12-qubits SVCs, since the amount of noise injected grows with the number of qubits, the collapse happens for smaller $p^{*}$.
-All methods seems to be more robust to amplitude damping error than to depolarizing and bit-flip errors.
-Figure 6: Accuracies (in percentage) of 6-qubits SVC as noise rate increases in the depolarizing, bit-flip and amplitude damping noise models.
+5 we plot, for each of the 3 methods, the accuracies achieved in the noiseless run and the 20 noisy runs.
+Figure 5: Accuracies of the quantum SVCs run on a noiseless simulator and on a noisy fake backend.
+In Table 4 we compare the performance of the noiseless models and the noisy models combined with the majority vote decision rule.
+Table 4: Accuracy (expressed in percentage) of quantum SVCs run on Aer Simulator without noise and on the noisy backend Fake Melbourne; for the latter the accuracy is obtained combining the predictions made on 20 executions with the majority vote decision rule.
+6q 8q 12q
+Aer Simulator 72.86 62.86 77.14
+Fake Melbourne 77.14 64.29 77.14
+The comparable or slightly better accuracy of the Fake Melbourne backend indicates that running multiple executions and using majority voting is a good strategy to mitigate variability caused by noise, achieving more consistent and accurate outcomes.
 4 Conclusion
-The aim of this work is comparing the performances of classical and quantum kernel methods for classifying samples of AD patients and healthy people subjected to handwriting tests.
-To achieve this, we have used first some classical approaches and then we have compared their performances with the ones achieved by quantum methods defined by various PQCs.
-Amongst classical SVCs, the one defined by the radial basis function kernel achieves average accuracy of 86.76%, while linear kernel and sigmoid kernel obtain 84.46% and 85.14% respectively.
-The quantum SVC defined by the 6-qubits PQC gets to 85.95%, the one defined by the 8-qubits PQC reaches 88.51% and the one defined by the 12-qubits circuit achieves average accuracy of 87.70%.
-The fact that the latter does not outperform the model defined by the 8-qubits circuit seems to indicate that performance does not automatically improve by increasing the number of qubits, and some ansatze are better suited for smaller PQCs, hence more feasible with the currently available quantum processors.
+The aim of this work is comparing the performances of classical methods and quantum kernel methods for classifying samples of AD patients and healthy people subjected to handwriting tests.
+The classical models used are Support Vector Classification, k-Nearest Neighbors and Decision Tree; quantum SVCs are defined by ansatze of 6, 8 and 12 qubits.
+Among the classical methods, the SVC delivers the highest performance, albeit with significant inconsistency: average accuracy is 85.28% and standard deviation 7.46%.
+Both kNN and DT are significantly less accurate.
+Evaluating the quantum methods we find that 6 and 8-qubits SVCs perform similarly to the classical SVC with slightly lower accuracy but better consistency.
+The quantum SVC with 12 qubits emerges as the most effective model, demonstrating superior accuracy and consistency.
+This indicates that increasing the number of qubits can significantly boost the performance of quantum classifiers, potentially unlocking more advanced capabilities and better results.
 In addition to running the methods considering all features together we have also defined methods based on data subsampling.
 We have considered separately the features associated with the 3 categories of tasks, namely features derived from graphic tasks, copy tasks or memory and dictation tasks.
-With graphic features, RBF and 8-qubits SVC achieve average accuracies of 79.19% and 79.86% respectively.
-When considering copy features, both RBF and 8-qubits SVC achieve average accuracy of 83.11%.
-Using the features derived from the tasks in the category ‘memory and dictation’, RBF and 8-qubits SVC achieve average accuracies of 73.65% and 71.89% respectively.
-Then we have examined the features obtained from each of the 25 task individually, defining a model for each task, and we have used the majority vote decision rule to combine the predictions made by the 25 models.
-Classical and quantum SVC obtain mean accuracies of 80.41% and 81.49% respectively.
-When restricted to only the best 5 tasks, i.e.
-the tasks that have achieved singularly the best accuracies in the previous runs, both classical and quantum SVC obtain average accuracies of 87.16%.
+In this case we just compare the performances of the 2 best models, i.e.
+classical SVC and 12-qubits SVC.
+The quantum algorithm generally shows improved or comparable performance across different feature subsets, highlighting its robustness and potential advantages in handling diverse types of data.
+Then we have examined the features obtained from each of the 25 task individually, defining a model for each task, and we have used the majority vote decision rule to combine the predictions made by the 25 models; after that we have considered just the best 5 tasks, i.e.
+the tasks that have achieved singularly the best accuracies in the previous runs.
+The quantum SVC demonstrates an advantage in accuracy over the classical SVC, whether considering all tasks or just the best ones, though it tends to have higher variability in its predictions.
 Kernels in quantum methods have been calculated simulating quantum computation without noise.
-We have also investigated how different types of noise affect the performance of quantum methods.
-From our analysis it appears that the methods show good robustness to the error introduced by amplitude damping, which is associated to energy dissipation.
-In fact, quantum SVCs tend to perform poorly only for high value of the noise rate related to the damping probability.
-Both depolarization and bit flip error instead cause the methods to perform worse for smaller values of noise rate.
-However, on real quantum computers different sources of noise act simultaneously, due to interaction with the environment and interaction between qubits, or imperfect application of quantum gates [29].
-Therefore quantum noise models, without in-depth physical experiments, can only partially tackle the complexity of this problem.
-Our study shows that quantum methods can achieve better performances in classification tasks even involving big datasets, as for the case of medical data analysis.
-Nowadays only QPUs with a limited number of noisy qubits are available.
-However this does not seem to affect too much the capability of quantum kernel methods.
-Moreover the constant improvement of quantum technologies suggests that there is a lot of room for enhancements to make quantum methods achieve even better results.
-The techniques used in our work may have a wide range of applications in QML algorithms for classification and regression, and can be implemented to enhance classical ML models for similar problems of early screening, e.g.
+We have also investigated how noise affects the performance of quantum methods using a fake backend, which mimics the behavior of a real noisy backend.
+Averaging the predictions from several executions on the noisy backend, with the majority rule, leads to a slightly superior accuracy, counteracting the variability introduced by noise.
+Several studies in classical machine learning have explored how introducing randomness and noise can enhance the performance of kernel methods [38][39].
+This concept holds significant promise when applied to quantum devices, which inherently operate in noisy environments.
+Noise may facilitate a broader exploration of the solution space, potentially leading to better solutions that might be missed in a noise-free environment.
+Regarding the use of real quantum hardware, currently, only IBM Quantum systems with 127 or more qubits are accessible, and each run for our computations takes 2 to 3 seconds.
+Similar considerations can be made for other quantum device providers.
+Despite computing multiple kernels in a single run, the execution of the methods exceeds the few minutes allocated in the free plan.
+Additionally, the number of kernels to compute increases quadratically with the number of samples.
+This makes quantum kernel methods more prone to time constraints compared to other quantum classifiers that can assign a label to a sample in a single run.
+Exploring optimizations, such as more efficient kernel computation techniques or hybrid quantum-classical approaches, could help mitigate these time constraints and make quantum kernel methods more feasible for real-world applications.
+While our experimental results are competitive with the current state of the art, there remains potential for further enhancement.
+In the quantum circuits we have used, only one hyperparameter, the bandwidth, is tuned in the validation phase.
+Enhancing the results could be achieved by incorporating one or more layers of parametric gates with tunable hyperparameters for optimization.
+Moreover, parameters ‘C’ , ‘gamma’ and ‘tol’ could be also added in the parameter grid of quantum methods.
+To enhance the user-friendliness of quantum methods, particularly in the clinical diagnostic process, we aim to automate the implementation of the methods we have developed.
+By simplifying the process of choice of the ansatz, optimization of parameters and execution on fake or real backends, we wish to make quantum kernel methods more accessible and usable for a wide range of applications, improving the performance of classical ML models for similar problems of early screening, e.g.
 detection of autism in children [40], or identification of various types of cancer [41, 42].